OpenStudy (anonymous):

A cup of coffee at 90 degrees Celsius is put into a 20 degree Celsius room when t=0 . The coffee’s temperature is changing at a rate of r(t)=-8(0.2^t) degree Celsius per minute, with t in minutes. Estimate the coffee’s temperature when t=10 . please helppp!medals are awarded!

OpenStudy (amistre64):

integrate with respect to t i think; since you know the rate of change perhaps?

OpenStudy (amistre64):

looks like a pain :)

OpenStudy (amistre64):

\[-8\int (.2^t)dt\]

OpenStudy (anonymous):

I'm not sure how to go about the limits of the integral though and I don't have the rate of change:x

OpenStudy (amistre64):

.2^t = y t = log.2(y) t = ln(y)/ln(.2) i c ant recall it too clearly either

OpenStudy (amistre64):

r(t)=-8(0.2^t) degree Celsius per minute, with t in minutes. Estimate the coffee’s temperature when t=10 r(t) = -8(.2^(10)) = -8.192 x10^-7

OpenStudy (amistre64):

http://www.wolframalpha.com/input/?i=-8%280.2^x%29

OpenStudy (amistre64):

so when t = 0; we got 90 degrees. 90 = 4.97068 * .2^t +C

OpenStudy (amistre64):

.2^0 = 1 :) C = 90 - 4.97068 = 85.02932 h(t) = 4.97068(2^10) + 85.02932 then

OpenStudy (amistre64):

i get: 85.0293.... for the temperature at 10 sec

OpenStudy (amistre64):

err.. 10 minutes

OpenStudy (anonymous):

you got that as the final? what were all the numerous steps in the link you sent me? oh my goodness, i appreciate this so much!

OpenStudy (amistre64):

the steps showed how they would have integrated it up to the indefinite; then you gotta pply your initial conditions to colve for the constant

OpenStudy (anonymous):

okay, I'm still looking it over now. I understand it for the most part. Thank you so so so so much! I am truly grateful for your help:)